# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

### v2

@@ 3.1 Graph-guided retrieval (G-retrieval)
In this section, we explore which forms of knowledge (e.g., triplets, paths, or subgraphs) should be extracted from KGs to enhance the LLMs and discuss the design of G-retrieval. The phase begins by encoding the user query into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$. For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are similarly computed using the same model, ensuring all the information is represented in a unified vector space. The relevance of each node or edge to the query is measured using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$. Based on these similarity scores, the most relevant nodes and edges are identified to form a set of top-$k$ nodes and edges for the query.
To balance retrieving as many relevant nodes and edges as possible while keeping the size manageable, we adopt a prize-cost trade-off strategy: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query. Specifically, we assign the top-$k$ nodes/edges with descending prize values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0. Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$. (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ during the expansion of the retrieved paths or subgraphs. These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
Triplet-retrieval: the triplet retrieval approach retrieves a fixed number $k$ of triplets with the highest total prize assigned to their respective head entity, relation, and tail entities.
Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fixed number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$. We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints. Finally, the top-scoring paths are sorted and returned. The details of path-retrieval can be found in Appendix A.4.
Sub-graph retrieval: the sub-graph retrieval approach also uses the prize assignment and cost allocation, following He et al. (2024a) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm. The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges. The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
@@ 3.2 Graph-guided generation (G-generator)
After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase learns to use this knowledge to generate the corresponding output to answer the user query. The generation is modeled as a sequential decision-making process, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)\vspace{-2mm}$ (2)
The model learns the parameters $\theta$ to maximize the likelihood of generating the correct output sequence by computing the probability distribution over possible next tokens at each time step. This ensures that the generated output not only considers the previously generated tokens but also the retrieved knowledge. The generation continues until a termination condition is met, such as generating an end-of-sequence token or reaching the maximum sequence length $T$.
@@ 4 Experiments
4.1 Evaluation Metrics
To quantify the hallucinations of LLMs, we leverage two public metrics, FActScore (Min et al., 2023) (factual precision in atomicity score) and SAFE (Wei et al., 2024) (search-augmented factuality evaluator). FActScore is designed to measure factual precision in text by decomposing a long form generation into multiple atomic facts and validates each separately against a reliable knowledge base, such as Wikipedia. We measure the proportion of facts that are supported by the knowledge source out of the total atomic facts. SAFE, on the other hand, adopts a more dynamic approach to fact-checking. It employs a language model as an investigative agent, which iteratively uses Google Search queries and reasons about whether the search results support or do not support the fact.
@@ 4.1 Evaluation Metrics
In addition to the hallucination metrics, we also propose four metrics to quantify the desirable qualities for generating sensible responses (Edge et al., 2024; Wang et al., 2023). Given the complexities inherent in an open-ended setting, collecting and annotating ground-truth answers poses significant challenges. Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm. To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics. The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question. Correctness: measures the clarity and specificity of the generated answer to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt can be found in Appendix C.
Models. We incorporate a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
@@ 4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings. We use zero-shot and few-shot prompting as baselines without injecting external knowledge. In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) as baselines. For LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 3) to study the impact of G-retrieval for reducing LLMs’ hallucinations. The results are shown in Table 2 and Figure 4. We obtain some intriguing findings:
Retrieving KG information can indeed mitigate factual errors in the responses. Methods integrating knowledge extracted from KGs show clear improvements in factual accuracy and comprehension scores compared to the baselines. For example, under Var-2 (triplet retrieval), GPT-4o achieves a FActScore of 72.55% $\pm$ 0.85%, which is a significant increase over the baseline score of 55.35% $\pm$0.95%. Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio. The radar chart in Figure 4 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection,” “Outcome Prediction,” and “Cause Explanation,” which require more reasoning and analysis rather than merely listing information. The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of the finding.
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can improve response quality (i.e., correctness, comprehensiveness, and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality. This shows that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
@@
Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries. We demonstrate the performance of different retrieval methods across different query types in Figure 4, showing that subgraphs achieve the best performance. Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning), subgraph retrieval methods achieve better performance compared to paths and triplets. Even for queries like “Relationship Explanation” and “Cause Explanation” which require stepwise reasoning, subgraph methods still demonstrate promising performance. This suggests that while different forms of retrieved knowledge offer unique benefits for specific types of queries, subgraphs provide consistently good performance.
@@ 4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5. It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings. Relation replacement has a relatively stronger impact on ATS, i.e., on the semantics of the perturbed KGs. For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs. We observe that (1) FS-SG consistently outperforms FS-T and FS-P even at higher perturbation levels, demonstrating its robustness by maintaining higher scores as perturbations increase. (2) FS-T and FS-P exhibit similar trends, each showing a significant performance drop as perturbation levels increase. Particularly, performance of FS-T and FS-P deteriorate when the perturbation level reaches 50%, i.e., becoming worse than the baseline using CoT. (3) On the setting using Relation Replacement which severely harms the semantics of the KGs, FS-T and FS-P decline more sharply than FS-SG. However, it still outperforms the baseline when the perturbation level is smaller than 40%.
Our findings on OKGQA-P illustrate that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance. We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT. In practical scenarios, platforms like Wikidata are less likely to be severely perturbed than 50% due to ongoing updates and community-based quality control, ensuring the relevance of our findings in real-world applications.
@@ 5 Conclusion
Appendix A Appendix
A.1 Intuition of Open-ended Questions?
In previous close-ended KGQA research Mihaylov et al. (2018); Yih et al. (2016); Talmor and Berant (2018), “hallucination” has not been considered at all. The primary reason is that close-ended KGQA models typically generate concise entity names as answers, which are directly compared against ground truth labels using metrics such as precision or accuracy. Any incorrect entity name is immediately reflected in the evaluation results, leaving little room for hallucination to manifest as a separate problem. While effective for simple tasks, we believe this closed-ended setting is too naive and fails to generalize to real-world, complex QA situations, where models must provide more detailed, nuanced responses Kantharaj et al. (2022).
In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers that include reasoning paths and supporting facts derived from the background KG. This broader output space not only provides a larger field to assess hallucination but also increases the likelihood of factual inaccuracies, as LLMs tend to produce more errors when generating longer and more complex responses. While hallucination is minimal in close-ended KGQA due to the short and factual nature of entity names, open-ended KGQA highlights the susceptibility of LLMs to hallucinate when generating rich, explanatory answers. By introducing this open-ended setting, OKGQA, we (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
A.2 Query Construction
In this section, we discuss the details of the query construction of OKGQA. It follows a template-based method with LLMs to generate a diverse range of queries. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7. Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as an automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assigned a set of human-label scores $s_{\text{human}}$ corresponding to $s_{\text{auto}}$, normalized to the same range as $s_{\text{auto}}$. We then optimize the input instructions by iteratively generating instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.

### v3

@@ 3.2 Graph-guided retrieval (G-retrieval)
Our goal in G-retrieval is to extract a compact yet informative subset $\mathcal{Z}^{*}$ from the KG that best supports answering the user query $q$. We first encode the query and all KG elements (nodes/edges) into a unified embedding space using a language model. We then measure the relevance of each element to $q$ (e.g., via cosine similarity) and identify a set of top-$k$ nodes and edges for the query.
To balance retrieving as many relevant nodes and edges as possible while keeping the $\mathcal{Z}^{*}$ size manageable, we adopt a prize-cost trade-off strategy Balas (1989) to guide the retrieval process: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query. Specifically, we assign the top-$k$ nodes/edges with descending prize values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0. Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$. (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ during the expansion of the retrieved paths or subgraphs. The final retrieval process aims to maximize the total prize (i.e., relevance) while minimizing associated costs.
We explore three retrieval variants for G-retrieval design (e.g., triplets, paths and subgraphs) as demonstrated in Figure 3.
  • Triplet-retrieval: retrieves a fixed number of triplets with the highest total prize assigned to their respective triplets.
  • Path-retrieval: starting from a fixed number of $k$ of high-prize nodes, we greedily expand paths $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ to maximize score: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$. We use a priority queue to iteratively return paths with top-scores and subject to maximum lengths and cycles. The details of path-retrieval can be found in Appendix A.3.
  • Sub-graph retrieval: building on He et al. (2024a), we use the Prize-Collecting Steiner Tree (PCST) algorithm to find a connected subgraph $\mathcal{S}$ that maximizes $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
@@ 3.3 Graph-guided Generation (G-Generator)
After retrieving $\mathcal{Z}^{*}$, the G-Generator use this knowledge to generate the paragraph-like response the user query. The generation is modeled as a sequential decision-making process: at each time step $t$, token $a_{t}$ is generated conditioned on $q$, $\mathcal{Z}^{*}$, and the previously generated tokens $a_{0:t-1}$:
$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}),$ (3)
where $\theta$ denotes the parameters of a neural text generation model. The generation stops when an end-of-sequence token is produced or when the maximum sequence length $T$ is reached.
@@ 4 Experiments
In this section, we first introduce the evaluation metrics, and then focus on two main research questions: RQ1: Can KGs reduce hallucination in LLMs? and RQ2: How are KG-Aware methods affected by noise/perturbations in KGs?
4.1 Evaluation Metrics & Setup
We quantify LLM hallucinations using two public metrics: FActScore (Min et al., 2023) and SAFE (Wei et al., 2024). FActScore measures factual precision by decomposing a long-form text into atomic facts and validating each against a reliable knowledge base like Wikipedia. In contrast, SAFE employs a language model as an investigative agent that iteratively employs Google Search queries to assess whether search results support the fact. For both metrics, we report the proportion of supported atomic facts out of the total atomic facts extracted from LLM responses.
In addition to the hallucination metrics, we propose four metrics using LLM-as-evaluator (Li et al., 2024) to quantify the quality of generated responses from LLM (Edge et al., 2024; Wang et al., 2023). In specific, we use G-Eval (Liu et al., 2023) framework for the evaluation and provide relevant Wikipedia pages of each query as context to enhance G-Eval’s robustness and stability. The four metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question. Correctness: measures the clarity and specificity of the generated answer to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt can be found in Appendix D.
We use gpt-4o-mini (from November 2024 to January 2025) as LLM backbone for all the evaluation metrics. As using LLM-as-evaluator frameworks may raise concern regarding potential self-enhancement or bias from the selection of the models Gu et al. (2024); Li et al. (2024), we conduct additional validation in Appendix A.4 (including human evaluation alignment and cross-validation across different LLM backbones), and find that the choice of LLM in the LLM-as-evaluator framework has little impact on the overall evaluation and demonstrate high correlation with the human evaluation, supporting the reliability of our testing.
For other testing LLM backbones mentioned in this section, we consider a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini (from November 2024 to January 2025), Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). Considering the trade-off between cost and performance, we use text-embedding-3-small model from OpenAI (from November 2024 to January 2025) as embedding model for G-retrieval process.
@@ 4.1 Evaluation Metrics & Setup
Baseline: With Wikipedia documents
GPT-4o - IRCoT $73.12\%\pm 0.32\%$ $69.23\%\pm 0.42\%$ $66.33\%\pm 0.34\%$ $65.51\%\pm 0.11\%$ $87.39\%\pm 0.68\%$ $69.45\%\pm 0.34\%$
GPT-4o-mini - IRCoT $70.31\%\pm 0.32\%$ $64.42\%\pm 1.31\%$ $61.37\%\pm 0.48\%$ $63.89\%\pm 0.72\%$ $84.72\%\pm 0.48\%$ $65.72\%\pm 1.03\%$
@@ 4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings. We use zero-shot and few-shot prompting as baselines without injecting external knowledge. In addition, we consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022), and more general RAG systems like IRCoT (Trivedi et al., 2022a) which retrieves paragraphs from Wikipedia to augment CoT generation. For LLMs augmented with KGs, we consider three KG retrieval variants: triplets, paths, and subgraphs to study the impact of G-retrieval for reducing LLMs’ hallucinations. The results are shown in Table 2 and Figure 4. We obtain some intriguing findings:
Retrieving KG information can indeed mitigate factual errors in the responses. Methods integrating knowledge extracted from KGs show clear improvements in factual accuracy and comprehension scores compared to the baselines. For example, under Var-2 (triplet retrieval), GPT-4o achieves a FActScore of 72.55% $\pm$ 0.85%, which is a significant increase over the baseline score of 55.35% $\pm$0.95%. Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio. The radar chart in Figure 4 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection,” “Outcome Prediction,” and “Cause Explanation,” which require more reasoning and analysis rather than merely listing information. The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of the finding. In addition, compared with RAG method IRCoT Trivedi et al. (2022b), leveraging Wikipedia documents, improves performance over zero-shot and 4-shot prompting by providing broad contextual support, it struggles with correctness and hallucination control due to the potential introduction of irrelevant or conflicting information. Our KG-based methods consistently outperform IRCoT, particularly in correctness, SAFE, and FActScore.
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like CoT and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can improve response quality (i.e., G-Eval) compared to baselines, but do not consistently improve factuality, and sometimes even diminish. This shows that relying solely on internal reasoning is inadequate for mitigating hallucinations, highlighting the necessity for external knowledge to address this issue effectively.
@@
Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries. We demonstrate the performance of different retrieval methods across different query types in Figure 4, showing that subgraphs achieve the best performance. Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning). Even for queries like “Relationship Explanation” and “Cause Explanation” which require stepwise reasoning, subgraph methods still demonstrate promising performance. This suggests that while different forms of retrieved knowledge offer unique benefits for specific types of queries, subgraphs provide consistently good performance.
@@ 4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5. It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings.
Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs. We observe that (1) FS-SG consistently outperforms FS-T and FS-P even at higher perturbation levels, demonstrating its robustness by maintaining higher scores as perturbations increase. (2) FS-T and FS-P exhibit similar trends, each showing a significant performance drop as perturbation levels increase. Particularly, performance of FS-T and FS-P deteriorate when the perturbation level reaches 50%, i.e., becoming worse than the baseline using CoT. (3) On the setting using Relation Replacement which severely harms the semantics of the KGs, FS-T and FS-P decline more sharply than FS-SG. However, it still outperforms the baseline when the perturbation level is smaller than 40%. In summary, we find that effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance. We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT. Considering practical scenarios, platforms like Wikidata are less likely to have perturbations as severe as 50% due to their ongoing updates and community-based quality control. This ensures the relevance and applicability of our findings in real-world settings.
@@ 6 Limitations
Appendix A Implementation Details
A.1 Query Construction
In this section, we discuss the details of the query construction of OKGQA. We first introduce the human-in-the-loop process to optimize the instruction for generating the queries, as shown in §A.1.1. We then present the metrics for quantify the generated queries in §A.1.2. Subsequently, we provide experiments results of human-in-the-loop process and demonstrate the Pearson correlation coefficients between human and LLM scores across rounds of optimization, and verify the inter-rater reliability across our evaluators in §A.1.3.
A.1.1 Human-in-the-loop for instruction optimization
To ensure that the generated queries accurately represent real-world scenarios and complexities, we propose a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7. To ensure clarity, we summarize this optimization process here:
  • Step 1: Generate a set of queries from an initial instruction.
  • Step 2: Collect automatic evaluation scores $s_{\mathbf{auto}}$ by LLMs and human-label scores $s_{\mathbf{human}}$ by human annotators for these queries (normalized to the same range).
  • Step 3: Identify patterns of discrepancies between these scores.
  • Step 4: Let the LLM analysis the identified patterns to generate new instructions,
The step 3 and 4 are conducted by prompting LLM with prompt specified in §D.3, and steps 1 to 4 are running iteratively to reducing $s_{\mathbf{auto}}$ and $s_{\mathbf{human}}$ discrepancies. This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.